### v1

@@ Abstract.
We extend the theoretical framework of proof mining by establishing general logical metatheorems that allow for the extraction of the computational content of theorems with prima facie “non-computational” proofs from probability theory, thereby unlocking a major branch of mathematics as a new area of application for these methods. Concretely, we devise proof-theoretically tame logical systems that, for one, allow for the formalization of proofs involving algebras of sets together with probability contents as well as associated Lebesgue integrals on them and which, for another, are amenable to proof-theoretic metatheorems in the style of proof mining that guarantee the extractability of effective and tame bounds from larges classes of ineffective existence proofs in probability theory. Moreover, these extractable bounds are guaranteed to be highly uniform in the sense that they will be independent of all parameters relating to the underlying probability space, particularly regarding events or measures of them. As such, these results, in particular, provide the first logical explanation for the success and the observed uniformities of the previous ad hoc case studies of proof mining in these areas and further illustrate their extent. Beyond these systems, we provide extensions for the proof-theoretically tame treatment of $\sigma$-algebras and associated probability measures using an intensional approach to infinite unions. Lastly, we establish a general proof-theoretic transfer principle that allows for the lift of quantitative information on a relationship between different modes of convergence for sequences of real numbers to sequences of random variables.
${}^{a}$ Department of Computer Science, University of Bath,
@@
${}^{b}$ Department of Mathematics, Technische Universität Darmstadt,
@@ 1. Introduction
One of the fundamental driving questions of proof theory is the following: What is the computational content of a mathematical theorem and how can it be exhibited? Proof mining, which emerged as a subfield of mathematical logic in the 1990s through the work of Ulrich Kohlenbach and his collaborators [^1] Historically, proof mining has its roots in Kreisel’s program of the “unwinding of proofs” [30, 31]. We refer to [23] for a comprehensive monograph on proof mining and its applications until 2008 and we refer to the survey article [28] for details on the earlier development of proof mining. aims at answering that question by extracting this computational content from theorems with proofs as they are found in the mainstream mathematical literature. This is, in particular, a non-trivial task as such proofs are prima facie noneffective, involving both classical logic as well as various non-computational (set-theoretic) principles. However, backed by a logical apparatus relying on the utilization of various methods from proof theory like functional interpretations and majorizability, this program of proof mining has had great success in various areas of mathematics, in particular regarding (nonlinear) analysis and optimization (see in particular the recent surveys [24, 25]).
Two areas where proof mining has previously only touched upon briefly are the fields of measure theory in general and probability theory in particular. Concretely, we refer to the works [1, 2, 3] which, so far, are essentially the only proof mining case studies in these areas. From a practical perspective, this diffidence of proof mining regarding these areas is at least partially due to the fact that they are so far not substantiated by underlying logical methods as other areas of applications for proof mining are, a fact that also renders the previous applications, to a certain degree, ad hoc.
@@
Concretely, the fundamental logical “substrate” of the proof mining program are the so-called general logical metatheorems. [^2] Examples of such logical metatheorems for proof mining can, e.g., be found in [8, 10, 13, 22, 23, 27, 34, 35, 41, 42, 43, 46]. These use well-known so-called proof interpretations like Gödel’s functional interpretation [11], negative translations (see e.g. [33]) and their extensions [^3] For example, other proof interpretations used in proof mining are Kreisel’s modified realizability interpretation [30] for the treatment of semi-constructive proofs, monotone variants of the Dialectica interpretation [20] to deal with the extraction of computable bounds from proofs or the related bounded functional interpretation [9]. to provide a general result that quantifies and allows for the extraction of the computational content of large classes of theorems from their proofs (which in particular may involve classical logic and various non-computational principles). In that way, proof mining, as substantiated by these metatheorems, has led to hundreds of applications in the last decades. A crucial innovation in the techniques underlying these logical metatheorems was introduced by Kohlenbach in [22], marking the “modern age” of proof mining: while the metatheorems for proof mining preceding [22] were based on “pure” systems for arithmetic in all finite types (see e.g. [23, 49]) and as such were restricted in their expressivity to dealing with spaces and structures that were representable as Polish metric spaces in the sense of Baire space (and thus separable), the paradigm first proposed in [22] was to extend the language of the underlying systems with additional abstract base types which, together with additional constants and governing axioms, could be used to talk about much larger and broader classes of spaces beyond merely representable ones. [^4] Whether a class of spaces can be treated in the context of the general logical metatheorems ultimately depends on the uniformity and complexity of the axioms describing said class as will also be discussed later in this paper. The class of spaces and objects treated in this fashion has grown since then to a rather sizable amount, ranging from fundamental examples like general (non-separable) metric, hyperbolic, $\mathrm{CAT}(0)$, Banach and Hilbert spaces to much more involved objects like $\mathbb{R}$-trees, $L^{p}$-spaces, the dual of a (non-separable) Banach space as well as monotone operators and nonlinear semigroups, among many others.
@@
Besides guaranteeing this high degree of uniformity of the extracted bounds, these metatheorems for probability theory further elucidate the extent of the phenomenon of so-called proof-theoretic tameness in the sense of Kohlenbach [26] (see also [37, 38] for related discussions of such phenomena), i.e. they show that also in the context of probability theory, one observes the empirical fact that most proofs, although in principle being subject to well-known Gödelian phenomena, nevertheless “seem to be tame in the sense of allowing for the extraction of bounds of rather low complexity”. Concretely, a crucial aspect of the approaches to various notions from probability theory taken here is that they avoid the computational strength inherent in some of the fundamental objects involved in this field (like the strength of comprehension needed to deal with countably infinite unions or Borel measurable functions, among other things) via the use of intensional methods and in that way, the complexity of the extractable information depends on (and can be a priori bounded in terms of) the complexity of the principles used in the corresponding proof and is not artificially high due to the abstract use of such notions. In that way, while the main base system taken in this paper is one that contains large amounts of comprehension (to illustrate the potential strength of systems which are amenable to proof mining methods), the approach taken here does not rely on these strong principles at all and thus can also immediately be developed over suitably weak subsystems (as e.g. the collection of systems introduced in [21] based on the Grzegorczyk hierarchy [12]) where most of the mathematics discussed here can be similarly carried out but where then bounds of correspondingly low complexity could be guaranteed and so the approach chosen here for these notions provides the right system to illustrate this apparent tameness of the area. [^5] Compare this in particular to the different approach by Kreuzer [32] towards extracting computational content from proofs in measure theory which, while being more limited in the scope of analyzable theorems, also in particular did not enjoy any of the uniformity or tameness features of the present approach.
As mentioned briefly before, the last two sections will further substantiate the applicability of these systems.
@@
For one, we outline in detail how the theorems analyzed in the seminal case study for proof mining in measure theory given in [2] formalize in our systems (and thereby, as mentioned before, we also provide the first logical explanation of the uniformities of the bounds extracted therein). In particular, connecting to the previous discussion on the theory of probability contents as a suitable base for a formal system for proof mining for general probability measures, the analysis of the results from [2] provided later shows that these remain true for probability contents and so, in particular, illustrate that the notions and proofs produced in the work [2] by following the finitary perspective of proof mining provide the right notions to simultaneously extend the analyzed results also to the theory of contents. In that way, this highlights the apparent empirical phenomenon that finitary quantitative variants of notions and results from the theory of probability measures, as suggested by proof mining, seem to provide suitable analogue notions for the theory of probability contents and it also highlights the naturalness of the theory of contents as an underlying medium for developing a logical account of probability theory in the sense of proof mining.
@@
Beyond that, we think that the present work lends itself both to further theoretical investigations on the extension of proof mining methods to further notions from probability and measure theory like Bochner integrals and martingales, among many others, as well as to substantiate and carry out many further (and potentially much more sophisticated) applications of proof mining in this area. In particular, we want to mention that most ideas developed here could be extended, mutatis mutandis, to general finite contents and measures.
@@ 2. Preliminaries
Lemma 2.1 (essentially [22, Lemma 2.10], see also [43, Lemma 2.1]).
@@ 3. Systems for algebras of sets
We write $\mathcal{F}^{\omega}$ for the system resulting from $\mathcal{A}^{\omega}$ over the augmented language including the types $\Omega,S$ (where all the respective constants and axioms now are allowed to also refer to these new types, if applicable) by extending this system with the constants $\mathrm{eq},\in,\cap,(\cdot)^{c},\emptyset,c_{\Omega}$ and the axioms $(\mathrm{eq})_{1}$$(\mathrm{eq})_{2}$ as well as $(\in)_{1}$$(\in)_{4}$.
@@ 4. Systems for contents on algebras of sets
    This follows via a simple induction from the axioms $(\mathbb{P})_{3}$.
@@ 5.1. Treating infinite unions tamely
Further, $\mathcal{F}^{\omega}[\bigcup]$ is closed under the following rule:
@@ 6.1. Intensional Intervals
The system $\mathcal{A}^{\omega}[\mathrm{Int}]$ proves the following properties of [,)normal-⋅normal-⋅[$\cdot$,$\cdot$)[ ⋅ , ⋅ ):
@@ 8. A bound extraction theorem
Lemma 8.6 ([10, 22], see also Kohlenbach [23, Lemma 17.80]).
@@ 9. Applications of the metatheorems
However, we want to mention that the focus on the work [2] shall be understood to be merely indicative on the usefulness of the systems and metatheorems discussed before and that essentially all other quantitative works on probability theory in the spirit of proof mining that have so far been considered in the literature and forthcoming work by the authors together with Thomas Powell can be similarly recognized as instances of the metatheorems proved here and, similarly, also there the peculiar uniformities observed in practice are a priori guaranteed by the approach towards the metatheorems chosen here.
@@
In particular, as this theorem of $\mathcal{F}^{\omega}[\mathbb{P}]$ has the correct logical form, our main Theorem 8.9 on the extraction of uniform computable bounds applies and we thus find that the existence of a computable bound on the existential quantifier on $n$ can be guaranteed to exist a priori and even further, based on our notion of majorizability, it can be guaranteed that this bound is independent of the content space and the sequence of events which matches exactly the properties of the bound explicitly calculuated in [2].
@@
and are done. So suppose $n\geq N$. Then by (*)(*)( * ), we get
@@ 10. Proof-theoretic transfer principles
$P_{0}(a,b,c,\widehat{x}):=\exists i^{0},j^{0}\left(b\leq_{0}i\land j\leq_{0}c\land|\widehat{x(i)}-\widehat{x(j)}|\in[0,2^{-a}]\right),$
@@
specifying that the properties $P_{0}$ and $Q_{0}$ (inducing the predicates $P$ and $Q$) induce measurable sets pointwisely relative to the sequence of random variables [^21] If we assume that $X$ is weakly Borel-measurable in the context of the above example (**)(**)( * * ), the corresponding point sets $P,Q$ as above indeed are measurable. specified by $X$ and that these random variables are all bounded via a suitable monotone sequence of bounds (i.e. that $X$ as a constant is majorized by $\tau$). It is clear that Theorem 8.9 extends to this system, which we denote by $\mathcal{U}^{\omega}$, as all constants are majorizable and since the new axioms are purely universal.
@@
If we consider the previous example for $P_{0}$ given in (**)(**)( * * ), then by formulating $P(X)$ a.u. in this case we recover the notion of almost uniform convergence with respect to finite unions as given in Definition 9.3 (i.e. the variant of almost uniform convergence implicitly considered in [2]).
@@
Acknowledgments: The first author was partially supported by the EPSRC Centre for Doctoral Training in Digital Entertainment (EP/L016540/1). The second author was supported by the ‘Deutsche Forschungsgemeinschaft’ Project DFG KO 1737/6-2. Both authors want to thank Ulrich Kohlenbach and Thomas Powell for insightful comments on various drafts on this paper.

### v2

@@ Abstract.
We extend the theoretical framework of proof mining by establishing general logical metatheorems that allow for the extraction of the computational content of theorems with prima facie “non-computational” proofs from probability theory, thereby unlocking a major branch of mathematics as a new area of application for these methods. Concretely, we devise proof-theoretically tame logical systems that, for one, allow for the formalization of proofs involving algebras of sets together with probability contents as well as associated Lebesgue integrals on them and which, for another, are amenable to proof-theoretic metatheorems in the style of proof mining that guarantee the extractability of effective and tame bounds from large classes of ineffective existence proofs in probability theory. Moreover, these extractable bounds are guaranteed to be highly uniform in the sense that they will be independent of all parameters relating to the underlying probability space, particularly regarding events or measures of them. As such, these results, in particular, provide the first logical explanation for the success and the observed uniformities of the previous ad hoc case studies of proof mining in these areas and further illustrate their extent. Beyond these systems, we provide extensions for the proof-theoretically tame treatment of $\sigma$-algebras and associated probability measures using an intensional approach to infinite unions. Lastly, we establish a general proof-theoretic transfer principle that allows for the lift of quantitative information on a relationship between different modes of convergence for sequences of real numbers to sequences of random variables.
a Department of Computer Science, University of Bath,
@@
b Department of Mathematics, Technische Universität Darmstadt,
@@ 1. Introduction
One of the fundamental driving questions of proof theory is the following: What is the computational content of a mathematical theorem and how can it be exhibited? Proof mining, which emerged as a subfield of mathematical logic in the 1990s through the work of Ulrich Kohlenbach and his collaborators, [^1] Historically, proof mining has its roots in Kreisel’s program of the “unwinding of proofs” [33, 34]. We refer to [26] for a comprehensive monograph on proof mining and its applications until 2008 and we refer to the survey article [31] for details on the earlier development of proof mining. aims at answering that question by extracting this computational content from theorems with proofs as they are found in the mainstream mathematical literature. This is, in particular, a non-trivial task as such proofs are prima facie noneffective, involving both classical logic as well as various non-computational (set-theoretic) principles. However, backed by a logical apparatus relying on the utilization of various methods from proof theory like functional interpretations and majorizability, this program of proof mining has had great success in various areas of mathematics, in particular regarding (nonlinear) analysis and optimization (see in particular the recent surveys [27, 28]).
Two areas where proof mining has previously only touched upon briefly are the fields of measure theory in general and probability theory in particular. Concretely, we refer to the works [1, 2, 3] which are essentially the only proof mining case studies in these areas preceding this work. From a practical perspective, this diffidence of proof mining regarding these areas is at least partially due to the fact that they are so far not substantiated by underlying logical methods as other areas of applications for proof mining are, a fact that also renders the previous applications, to a certain degree, ad hoc.
@@
Concretely, the fundamental logical “substrate” of the proof mining program are the so-called general logical metatheorems. [^2] Examples of such logical metatheorems for proof mining can, e.g., be found in [10, 12, 16, 25, 26, 30, 37, 38, 46, 47, 48, 49, 52]. These use well-known so-called proof interpretations like Gödel’s functional interpretation [13], negative translations (see e.g. [36]) and their extensions [^3] For example, other proof interpretations used in proof mining are Kreisel’s modified realizability interpretation [33] for the treatment of semi-constructive proofs, monotone variants of the Dialectica interpretation [23] to deal with the extraction of computable bounds from proofs or the related bounded functional interpretation [11]. to provide a general result that quantifies and allows for the extraction of the computational content of large classes of theorems from their proofs (which in particular may involve classical logic and various non-computational principles). In that way, proof mining, as substantiated by these metatheorems, has led to hundreds of applications in the last decades. A crucial innovation in the techniques underlying these logical metatheorems was introduced by Kohlenbach in [25], marking the “modern age” of proof mining: while the metatheorems for proof mining preceding [25] were based on “pure” systems for arithmetic in all finite types (see e.g. [26, 56]) and as such were restricted in their expressivity to dealing with spaces and structures that were representable as Polish metric spaces in the sense of Baire space (and thus separable), the paradigm first proposed in [25] was to extend the language of the underlying systems with additional abstract base types which, together with additional constants and governing axioms, could be used to talk about much larger and broader classes of spaces beyond merely representable ones. [^4] Whether a class of spaces can be treated in the context of the general logical metatheorems ultimately depends on the uniformity and complexity of the axioms describing said class as will also be discussed later in this paper. The class of spaces and objects treated in this fashion has grown since then to a rather sizable amount, ranging from fundamental examples like general (non-separable) metric, hyperbolic, $\mathrm{CAT}(0)$, Banach and Hilbert spaces to much more involved objects like $\mathbb{R}$-trees, $L^{p}$-spaces, the dual of a (non-separable) Banach space as well as monotone operators and nonlinear semigroups, among many others.
@@
Besides guaranteeing this high degree of uniformity of the extracted bounds, these metatheorems for probability theory further elucidate the extent of the phenomenon of so-called proof-theoretic tameness in the sense of Kohlenbach [29] (see also [40, 41] for related discussions of such phenomena), i.e. they show that also in the context of probability theory, one observes the empirical fact that most proofs, although in principle being subject to well-known Gödelian phenomena, nevertheless “seem to be tame in the sense of allowing for the extraction of bounds of rather low complexity”. Concretely, a crucial aspect of the approaches to various notions from probability theory taken here is that they avoid the computational strength inherent in some of the fundamental objects involved in this field (like the strength of comprehension needed to deal with countably infinite unions or Borel measurable functions, among other things) via the use of intensional methods and in that way, the complexity of the extractable information depends on (and can be a priori bounded in terms of) the complexity of the principles used in the corresponding proof and is not artificially high due to the abstract use of such notions. In that way, while the main base system taken in this paper is one that contains large amounts of comprehension (to illustrate the potential strength of systems which are amenable to proof mining methods), the approach to the various objects from probability theory taken here does not rely on these strong principles at all and thus can also immediately be developed over suitably weak subsystems (as e.g. the collection of systems introduced in [24] based on the Grzegorczyk hierarchy [15]) where most of the mathematics discussed here can be similarly carried out but where then bounds of correspondingly low complexity could be guaranteed and so the approach chosen here for these notions provides the right system to illustrate this apparent tameness of the area.
The present approach to proof mining and probability theory is in particular to be distinguished with previous work on logical aspects of (quantitative) probability theory. On the proof-theoretic side, the main preceding work is that by Kreuzer [35] on extracting computational content from proofs in measure theory which however relies on strong forms of comprehension for treating measure spaces, resulting in a rather restricted formal theory with a more limited scope of analyzable theorems, and which further does not guarantee any of the uniformity features for the quantitative information extracted thereby in contrast to the present approach. Outside of proof theory, finitizations of concepts from probability theory that enjoy similar uniformities as the ones considered here have also been obtained using tools from model theory, particularly ultraproducts [4, 8, 14], though differing from the results presented here in both method and scope. The first crucial difference is that the proof-theoretic metatheorems presented here actually provide a method for extracting the uniform quantitative information of which the model-theoretic approach only can infer the existence of. Beyond that, as already commented on before, in the present approach the complexity of that information can be gauged beforehand based on the principles used in the proof and the proof-theoretic tameness of that information can be guaranteed thereby. Lastly, the model-theoretic approach is essentially fixed to focus on convergence statements and uniformities relating to their so-called metastable formulations whereas the proof-theoretic approach presented here does not have that limitation (which has a crucial impact on the range of possible applications as will also be discussed again in the end of the introduction). These facts crucially separate our work from the model-theoretic approach. Nevertheless, the model-theoretic approaches are certainly not subsumed by our work but rather complementary. For one, the model-theoretic approaches to uniformities of rates of metastability in probability theory only rely on the truth of a statement while our proof-theoretic results rely on the provability of the underlying theorem in an (albeit very strong) underlying theory. Also, we want to highlight that in particular the approach by Dueñez and Iovino [8] seems to rely, upon a closer inspection, on rather similar ideas for approaching some of the initial objects in question (e.g. by effectively treating probability contents on algebras instead of measures on $\sigma$-algebras and treating these underlying spaces with their operations as abstract entities), a fact that illustrates the apparent similarity of the problems that both the proof-theoretic and the model-theoretic approaches to these types of questions face and which in particular further highlights the naturalness of the approach followed here. However, our approach starts to crucially differ also conceptually instead of just methodologically from the work in [8] in the way that infinite unions and integration are treated in our systems and in the way that the uniformities of associated (extracted) bounds are guaranteed.
Still, besides these logical considerations it is the applications of the present metatheorems to probability theory which are arguably the most important consequence of the present work and we hence use the last two sections of this paper to substantiate the applicability of the logical metatheorems introduced here.
@@
For one, we outline in detail how the quantitative analysis of Egorov’s theorem as presented in the seminal case study for proof mining in measure theory from [2] formalizes in our systems (and thereby, as mentioned before, we also provide the first logical explanation of the uniformities of the bounds extracted therein [^5] This is in particular to be compared with the work by Dueñez and Iovino [8] where a model-theoretic account of the uniformities of a quantiative variant of the dominated convergence theorem, partially akin to the results of [2], is given. These considerations of [8] however avoid Egorov’s theorem and with that the central object of study in [2] by which their quantitative version of the dominated convergence theorem is established therein.). In particular, connecting to the previous discussion on the theory of probability contents as a suitable base for a formal system for proof mining for general probability measures, the analysis of the results from [2] provided later highlights that these remain true for probability contents and so, in particular, illustrate that the notions and proofs produced in the work [2] by following the finitary perspective of proof mining provide the right notions to simultaneously extend the analyzed results also to the theory of contents. In that way, this points to the apparent empirical phenomenon that finitary quantitative variants of notions and results from the theory of probability measures, as suggested by proof mining, seem to provide suitable analogue notions for the theory of probability contents and it also highlights the naturalness of the theory of contents as an underlying medium for developing a logical account of probability theory in the sense of proof mining. In that way, the fundamental relevance of (probability) contents is essentially rediscovered by the proof-theoretic approach presented in this paper.
@@
Besides these examples of application discussed here, the applicability of the systems presented in this paper is further substantiated by the fact the they explain the previously mentioned application [1] preceding this work as well as the very recent works [43, 44] (together with forthcoming work by the authors together with Thomas Powell) as instances of the present methodology. In particular, in the context of the works [43, 44], the extractive proof-theoretic perspective of this work was crucial for obtaining the respective results.
Beyond that, we think that the present work lends itself both to further theoretical investigations on the extension of proof mining methods to further notions from probability and measure theory like Bochner integrals and martingales, among many others, as well as to substantiate and carry out many further (and potentially much more sophisticated) applications of proof mining in this area beyond the already mentioned. In particular, we want to mention that most ideas developed here could be extended, mutatis mutandis, to general finite contents and measures.
@@ 2. Preliminaries
Lemma 2.1 (essentially [25, Lemma 2.10], see also [49, Lemma 2.1]).
@@ 3. Systems for algebras of sets
We write $\mathcal{F}^{\omega}$ for the system resulting from $\mathcal{A}^{\omega}$ over the augmented language including the types $\Omega,S$ (where all the respective constants and axioms now are allowed to also refer to these new types, if applicable) by extending this system with the constants $\mathrm{eq},\in,\cup,(\cdot)^{c},\emptyset,c_{\Omega}$ and the axioms $(\mathrm{eq})_{1}$$(\mathrm{eq})_{2}$ as well as $(\in)_{1}$$(\in)_{4}$.
@@
Remark 3.4.
Also the constants $\mathrm{eq}$ and $\in$ are immediately provably extensional in $\mathcal{F}^{\omega}$, as can be shown using the quantifier-free extensionality rule.
@@ 4. Systems for contents on algebras of sets
    This follows via a simple induction from the axiom $(\mathbb{P})_{3}$.
@@ 5.1. Treating infinite unions tamely
Further, in $\mathcal{F}^{\omega}[\bigcup]$ the following rule is derivable:
@@ 6.1. Intensional Intervals
The system $\mathcal{A}^{\omega}[\mathrm{Int}]$ proves the following properties of [,)[$\cdot$,$\cdot$)[ ⋅ , ⋅ ):
@@ 8. A bound extraction theorem
Lemma 8.6 ([12, 25], see also Kohlenbach [26, Lemma 17.80]).
@@ 9. Applications of the metatheorems
However, as already discussed in the introduction, we want to mention that the focus on the work [2] shall be understood to be merely indicative on the usefulness of the systems and metatheorems discussed before and that essentially all other quantitative works on probability theory in the spirit of proof mining that have so far been considered in the literature (as well as forthcoming work by the authors together with Thomas Powell) can be similarly recognized as instances of the metatheorems proved here and, similarly, also there the peculiar uniformities observed in practice are a priori guaranteed by the approach towards the metatheorems chosen here. In particular, as discussed in the introduction, the logical perspective provided by this work was crucial for obtaining the recent case studies presented in [43, 44].
@@
In particular, as this theorem of $\mathcal{F}^{\omega}[\mathbb{P}]$ has the correct logical form, our main Theorem 8.9 on the extraction of uniform computable bounds applies and we thus find that the existence of a computable bound on the existential quantifier on $n$ can be guaranteed to exist a priori and even further, based on our notion of majorizability, it can be guaranteed that this bound is independent of the content space and the sequence of events which matches exactly the properties of the bound explicitly calculated in [2].
@@
and are done. So suppose $n\geq N$. Then by ()(*)( ∗ ), we get
@@ 10. Proof-theoretic transfer principles
$P_{0}(a,b,c,\widehat{x}):=\forall i^{0},j^{0}\left(b\leq_{0}i,j\land i,j\leq_{0}c\to|\widehat{x(i)}-\widehat{x(j)}|\in[0,2^{-a}]\right),$
@@
specifying that the properties $P_{0}$ and $Q_{0}$ (inducing the predicates $P$ and $Q$) induce measurable sets pointwisely relative to the sequence of random variables [^21] If we assume that $X$ is weakly Borel-measurable in the context of the above example ()(**)( ∗ ∗ ), the corresponding point sets $P,Q$ as above indeed are measurable. specified by $X$ and that these random variables are all bounded via a suitable monotone sequence of bounds (i.e. that $X$ as a constant is majorized by $\tau$). It is clear that Theorem 8.9 extends to this system, which we denote by $\mathcal{U}^{\omega}$, as all constants are majorizable and since the new axioms are purely universal.
@@
If we consider the previous example for $P_{0}$ given in ()(**)( ∗ ∗ ), then by formulating $P(X)$ a.u. in this case we recover the notion of almost uniform convergence with respect to finite unions as given in Definition 9.3 (i.e. the variant of almost uniform convergence implicitly considered in [2]).
@@
Acknowledgments: The first author was partially supported by the EPSRC Centre for Doctoral Training in Digital Entertainment (EP/L016540/1). The second author was supported by the ‘Deutsche Forschungsgemeinschaft’ Project DFG KO 1737/6-2. Both authors want to thank Thomas Powell, Ulrich Kohlenbach, Henry Towsner and José Iovino for insightful discussions on the topics of this paper.
Declarations of interest: none